# A Gaussian Sliding Windows Regression Model for Hydrological Inference

### v3

@@ 1 Introduction
Hence we investigate the problem of inferring streamflow partitioning into different flow paths such as overland flow, subsurface flow, and baseflow (Cornette et al., 2022; Kannan et al., 2007; McMillan, 2020; Nejadhashemi et al., 2009; Wang and Gupta, 2024). While there exist many simple methods which can partition streamflow into fast surface flow versus slow baseflow (Nejadhashemi et al., 2009; Wang and Gupta, 2024), the processes and reasoning behind this partitioning are usually lost. To prevent this, one could consider how input rainfall is partitioned into different flow paths. Hence, we consider a time series $\bm{y}_{T}=\left(y_{t}\right)_{t\in T}$ of streamflow at a particular gauge and an input time series $\bm{x}_{T}=\left(x_{t}\right)_{t\in T}$ of rainfall data on a common observational interval $T$. Given such data, a common approach is a lagged regression model following the concepts of auto-regressive moving-average with exogenous variables (ARIMAX, ARMAX) (Hyndman and Athanasopoulos, 2021). In contrast to these models, our problem does not involve lagged observations of the target $y_{t}$; instead, we seek to fully characterize the target time series via the lagged input time series. Further, we desire more interpretability and inferential power than is offered by ARIMAX models. For example, the maximum time interval for which the predictor has an impact on the response and the separation of flow pathways is of interest for watershed modeling.
A more closely related model in time series analysis is the finite distributed lag model (DLM) (Baltagi, 2022). Despite its roots in econometrics, the DLM has been extensively applied in environmental sciences (Chen et al., 2018; Peng et al., 2009; Rushworth et al., 2013; Warren et al., 2020). While ARIMAX mainly relies on the concept of ARIMA, i.e. it makes use of the lagged target variable $y_{t}$ as input, and only adds the present time point of the exogenous variable $x_{t}$, DLM predicts the target variable purely based on a distinct lagged input time series — the same setup that is pursued in this work. In this sense, the presented work may be seen as an extension of the work by Rushworth et al. (2013), however, with an improved parameterization, which allows for improved model interpretability. Basic DLM variants do not allow us to impose parametric assumptions on the shape of the regression parameter vector beyond simple singular shapes like geometrically decreasing patterns or polynomials (Almon, 1965; Eisner, 1960; Griliches, 1967), though recent work (Rushworth, 2018) extends the concept to B-splines in a Bayesian setting. Such DLM extensions permit a large variety of lag curve shapes, but do not make the separation between flow paths explicit. Crucially, then, they do not allow users to extract hydrologic inferences about distinct flow pulses.
@@ 2 Sliding Windows Regression Model
Our proposed model pursues the idea that the effect of precipitation on streamflow follows a mixture of $k$ kernels associated with temporal windows, each one characterizing a particular flow path (groundwater flow, sub-surface flow, or overland flow). Ultimately, we aim to represent the distribution of response times from flow path $i\in\{1,\dots,k\}$ by weight vector $\bm{\kappa}^{(i)}\in\mathbb{R}^{t+1}$. Computing a weighted sum of the lagged time series $\bm{x}_{[t]}$ with one such weight vector $\bm{\kappa}^{(i)}$ models the contribution to streamflow $y_{t}$ associated with flow path $i$. Summing the contributions gives the overall flow. To facilitate explanations, we present the model for the special case of only $k=1$ window first, and extend it to multi-window scenarios afterwards.
@@ 2.1 Single-window model
Figure 2: Illustration of the Sliding Windows Regression model with $k=1$ window, which predicts $y_{t}=y_{70}$ based on a subset of ${\bm{x}}_{[t]}$ values in a window. With $\delta=20$ and $r=10$, the window $[40,60]$ centred around $t-\delta=50$ covers $212r+1=$ 21 time points.
@@ 2.2 Multi-window model
Assuming multiple different (potentially overlapping) windows allows the model to account for distinct effects on the current target $y_{t}$. For this purpose, we extend our method from a single-window model to a mixture of $k$ kernels, which is related to the density resulting from a mixture of probability distributions (Everitt et al., 1981). For a given window $i\in\{1,\dots,k\}$, we use the same notation as in the single-window case for the window and kernel parameters with an added index $i$ denoting the window number: $\delta_{i}$, $r_{i}$, and $\sigma_{i}$. The Gaussian SWR model with multiple windows is given as a direct generalization of (1) by
$y_{t}$ $=\sum_{i=1}^{k}\beta_{i}\cdot\left({\bm{x}}_{[t]}\ast\bm{\kappa}_{i}\right)+\varepsilon_{t},$ (3)
$={\bm{x}}_{[t]}\ast\left(\sum\limits_{i=1}^{k}\beta_{i}\bm{\kappa}_{i}\right)+\varepsilon_{t}$ (4)
where the regression parameters $\beta_{1},\dots,\beta_{k}\geq 0$ act as window weights. In summary, the Gaussian SWR model with $k$ windows contains the following parameters
  • $k$ regression parameters ($\beta_{1},\dots,\beta_{k}$), $\beta_{i}\in\mathbb{R}^{+}$,
  • $k$ lag parameters ($\delta_{1},\dots,\delta_{k}$), $\delta_{i}\in\mathbb{R}^{+}$, and
  • $k$ kernel variance (size) parameters ($\sigma_{1}^{2},\dots,\sigma_{k}^{2}$), $\sigma_{i}\in\mathbb{R}^{+}$.
The concept of a Gaussian SWR model with multiple windows is illustrated in Fig. 4: a target value $y_{t}$ is predicted by multiple windows accumulating information in ${\bm{x}}_{[t]}$ over different time intervals. Hence, the joint kernel forms a multi-modal distribution and can be interpreted as a kernel density estimator describing a more general mixture distribution of Gaussian and truncated Gaussian kernels. As depicted in Fig. 5, the combined kernel represents a linear combination of the single window kernels $\bm{\kappa}_{i}$ for $i=1,\dots,k$. Under the assumption that each window has a distinct centre point $\delta_{i}$ (otherwise, identifiability of the parameters might not be given), without loss of generality we can order the location parameters such that $\delta_{1}<\delta_{2}<\dots<\delta_{k}$.
@@
Figure 4: Illustration of the Sliding Windows Regression model with $k=2$ windows, which predicts $y_{t}=y_{70}$ based on $x_{[70]}$. The window location parameters are $\delta_{1}=20$ and $\delta_{2}=40$. On the time axis, the window $[40,60]$ is centred around $t-\delta_{1}=50$ and covers $r_{1}=10$ time points on each side of $\delta_{1}=50$, while window $[10,50]$ is centred around $t-\delta_{2}=30$ and covers $r_{2}=20$ time points on each side.
@@ 2.2.1 Parameter interpretation
By definition, the number of windows $k$ models the number of underlying, separable flow paths.
Since single-window kernels $\bm{\kappa}$ are normalized, regression parameters (window coefficients) $\beta_{1},\dots,\beta_{k}$ act as weights among the kernels. For practical evaluations, they may be converted into proportions representing the relative impacts of the underlying flow paths.
For a window that is essentially untruncated, the location parameter $\delta_{i}$ and the width $\sigma_{i}$ of kernel $i$ indicate the mean and the standard deviation, respectively, of the time lag between precipitation and a change in streamflow. In any case $\delta_{i}$ is always the window mode and is thus a direct estimate of the typical time delay between rainfall events and the resulting impact on the gauged runoff for a flow path.
@@
Figure 5: Example mixture of window kernels (weights): kernel $\bm{\kappa}_{1}$ represents a truncated Gaussian, while kernel $\bm{\kappa}_{2}$ has a Gaussian shape. The combined window $\beta_{1}\bm{\kappa}_{1}+\beta_{2}\bm{\kappa}_{2}$ is a mixture of Gaussians for the SWR model.
@@ 2.3 Information criteria
$\varepsilon_{t}=y_{t}-{\bm{x}}_{[t]}\ast\left(\sum\limits_{i=1}^{k}\beta_{i}\bm{\kappa}_{i}\right),$
@@ 2.4 Model training & implementation
    $\bm{D}^{(0)}$ (a vector of starting values for the $\delta_{i}$)
    $\bm{S}^{(0)}$ (a vector of starting values for the $\sigma_{i}$)
@@
         $\bm{B}^{(1)}\leftarrow(m_{y}/m_{x},\ldots,m_{y}/m_{x})^{\top}$triangleright (for $\beta_{1}$)
         $\bm{D}^{(1)}\leftarrow(\bm{D}^{(0)},\bm{D}^{(0)},\ldots)^{\top}$triangleright (for $\delta_{1}$)
         $\bm{S}^{(1)}\ \leftarrow(\bm{S}^{(0)},\bm{S}^{(0)},\ldots)^{\top}$triangleright (for $\sigma_{1}$)
@@
         Generate an $n\times 1$ vector of starting values for $\beta_{1},\ldots,\beta_{k}$
@@
                  $\bm{B}^{(j)}\leftarrow(m_{y}/m_{x}/k,\ldots,m_{y}/m_{x}/k)$triangleright (for $\beta_{j}$)
@@
         Generate an $n\times 1$ vector of starting values for $\delta_{k}$ (note $\bm{D}^{(1)},\ldots,\bm{D}^{(k-1)}$ are unchanged)
@@
         Generate an $n\times 1$ vector of starting values for $\sigma_{k}$ (note $\bm{S}^{(1)},\ldots,\bm{S}^{(k-1)}$ are unchanged)
@@
     Generate an $n\times(3k)$ population of starting values with columns for $\beta_{1},\ldots,\beta_{k},\delta_{1},\ldots,\delta_{k},\sigma_{1},\ldots,\sigma_{k}$
@@
We use the GENOUD algorithm (Mebane and Sekhon, 2011) in the R package rgenoud (Mebane, W. R., Sekhon, J. S., Saarinen, T., 2023) to train a model with a fixed number of windows. GENOUD optimizes a population of candidate solutions over a number of generations via operations analogous to genetic evolution. It also fine tunes the best candidate found at each generation with a quasi-Newton method. Trial and error comparisons with a more standard constrained optimizer (BOBYQA due to Powell, 2009) demonstrated a clear advantage in the optimal log likelihood found, which is important if BIC is to reliably choose the number of windows. As a stopping criterion, the absolute tolerance is set to $10^{-3}$, which is a trivial change in the log likelihood.
@@
In the GENOUD algorithm, we constrain the parameters such that all $\delta_{i}>0$ and all $\sigma_{i}>1/6$, which allows for truncated as well as non-truncated kernels. In the boundary case of $\delta=0$, a maximum level of truncation is reached for the Gaussian kernel. The constraint $\sigma_{i}>1/6$, avoids collapse of the window. Moreover, the minimum window width is hence 1, i.e., ±$\pm 3\times 1/6$, which is commensurate with the limitation of discrete data at unit time steps.
@@
GENOUD requires a population of $n$ starting values for $\beta$, $\delta$, and $\sigma$. Experimental observations showed that this initial population is important. It needs to provide coverage of reasonable values in the parameter space but not be so overly wide as to make the optimization inefficient. Note that the initial ranges can be exceeded later during the GENOUD iterations. The population formed in Alg. 1 has $n$ members in the rows of a matrix ($n=100$ for all results reported) and $3k$ columns for initial values of $\beta_{1},\ldots,\beta_{k},\delta_{1},\ldots,\delta_{k},\sigma_{1},\ldots,\sigma_{k}$, respectively. For $k=1$, $\beta_{1}$ is always set to $m_{y}/m_{x}$, where $m_{x}$ and $m_{y}$ are the sample means of $\bm{x}_{T}$ and $\bm{y}_{T}$, respectively, to respect the scaling of the data when a single kernel has total unit weight. The $\delta_{1}$ starting values repeat a vector $\bm{D}^{(0)}$; we use $\bm{D}^{(0)}=(1,2,\ldots,25)^{T}$, which will be repeated four times to fill the population of size $n=100$. This range is reasonable for the hydrology application but would be changed in other contexts. Similarly, the $\sigma_{1}$ starting values repeat $\bm{S}^{(0)}=(1,5,10,20)^{T}$. When $k>1$ and window $k$ is added, $\beta_{1},\ldots,\beta_{k}$ are all reduced in magnitude to $m_{y}/m_{x}/k$ to account for $k$ kernels. The starting values for $\delta_{1}$ and $\sigma_{1}$ are as for the $k=1$ iteration; the $\delta_{k}$ and $\sigma_{k}$ for the further windows are given by a random permutation of the elements in $\bm{D}^{(0)}$ and $\bm{S}^{(0)}$, respectively. This procedure ensures the starting values for a multi-window iteration have sufficient coverage of the parameter space, while having distinct initial values for each of the $k$ windows.
@@
The Gaussian SWR model is implemented in R version 4.3.0 (R Core Team, 2022). The implementation is publicly available on GitHub [^1] https://www.github.com/sschrunner/SlidingWindowReg.

### v4

@@ 1 Introduction
Hence we investigate the problem of inferring streamflow partitioning into different flow paths such as overland flow, subsurface flow, and baseflow (Cornette et al., 2022; Kannan et al., 2007; McMillan, 2020; Nejadhashemi et al., 2009; Wang and Gupta, 2024). While there exist many simple methods which can partition streamflow into fast surface flow versus slow baseflow (Nejadhashemi et al., 2009; Wang and Gupta, 2024), the processes and reasoning behind this partitioning are usually lost. To prevent this, one could consider how input rainfall is partitioned into different flow paths. Hence, we consider a time series $\bm{y}_{T}=\left(y_{t}\right)_{t\in T}$ of streamflow at a particular gauge and an input time series $\bm{x}_{T}=\left(x_{t}\right)_{t\in T}$ of rainfall data on a common time domain $T$. Given such data, a common approach is a lagged regression model following the concepts of auto-regressive moving-average with exogenous variables (ARIMAX, ARMAX) (Hyndman and Athanasopoulos, 2021). In contrast to these models, our problem does not involve an auto-regressive part, i.e., we do not use lagged observations of the target $y_{t}$ as predictors; instead, we seek to fully characterize the target time series via the lagged input time series. Further, we desire more interpretability and inferential power than is offered by ARIMAX models. For example, the maximum time interval for which the predictor has an impact on the response and the separation of flow pathways is of interest for watershed modeling.
A more closely related model in time series analysis is the finite distributed lag model (DLM) (Baltagi, 2022). Despite its roots in econometrics, the DLM has been extensively applied in environmental sciences (Chen et al., 2018; Peng et al., 2009; Rushworth et al., 2013; Warren et al., 2020). While ARIMAX mainly relies on the concept of ARIMA, i.e. it makes use of the lagged target variable $y_{t}$ as input, and only adds the present time point of the exogenous variable $x_{t}$, DLM predicts the target variable purely based on a distinct lagged input time series — the same setup that is pursued in this work. In this sense, the presented work may be seen as specific type of a DLM and thereby an extension of the work by Rushworth et al. (2013), however, with an improved parameterization, which allows for improved model interpretability. Basic DLM variants do not allow us to impose parametric assumptions on the shape of the regression parameter vector beyond simple singular shapes like geometrically decreasing patterns or polynomials (Almon, 1965; Eisner, 1960; Griliches, 1967), though recent work (Rushworth, 2018) extends the concept to B-splines in a Bayesian setting. Such DLM extensions permit a large variety of lag curve shapes, but do not make the separation between flow paths explicit. To the best of our knowledge, none of the existing DLM variants is capable of explicitly separating distinct flow pulses, which is most crucial for hydrologic inference.
@@ 2 Sliding Windows Regression Model
Our proposed model pursues the idea that the effect of precipitation on streamflow follows a mixture of $k$ kernels associated with temporal windows, each one characterizing a particular flow path (groundwater flow, sub-surface flow, or overland flow). Ultimately, we aim to represent the distribution of response times from flow path $i\in\{1,\dots,k\}$ by weight vector $\bm{\kappa}^{(i)}\in\mathbb{R}^{t+1}$. Note that the length of $\bm{\kappa}^{(i)}$ is formally time-dependent based on its dimension. Since the sub-vector of non-zero weights remains the same over time, zero padding is used at the end to comply with the dimensionality of $\bm{x}_{[t]}$. Thus, the time index is left out in a slight abuse of notation. Computing a weighted sum of the lagged time series $\bm{x}_{[t]}$ with one such weight vector $\bm{\kappa}^{(i)}$ models the contribution to streamflow $y_{t}$ associated with flow path $i$. Summing the contributions gives the overall flow. To facilitate explanations, we present the model for the special case of only $k=1$ window first, and extend it to multi-window scenarios afterwards.
@@ 2.1 Single-window model
Unlike typical linear regression models, the SWR model is defined without an intercept $\beta^{(0)}$. This is due to the assumption that runoff $y_{t}$ is 00 after a long period without precipitation, $x_{s}=0$ for all $s<t$. However, in other applications, the proposed SWR model may be used with an intercept.
Further, note that the model definition allows negative responses $y_{t}$ if all entries of $\beta$ are very small while the error $\varepsilon_{t}$ takes a larger negative value. However, in this rare case, point forecasts remain non-negative due to non-negative observations $\bm{x}_{[t]}$.
@@
Figure 2: Illustration of the Sliding Windows Regression model with $k=1$ window, which predicts $y_{70}$ based on a subset of ${\bm{x}}_{[70]}$ values in a window. With $\delta=20$ and $r=10$, the window $[40,60]$ centred around $s-\delta=50$ covers $212r+1=$ 21 time points.
@@ 2.2 Multi-window model
Assuming multiple different (potentially overlapping) windows allows the model to account for distinct effects on the current target $y_{t}$. For this purpose, we extend our method from a single-window model to a mixture of $k$ kernels, which is related to the density resulting from a mixture of probability distributions (Everitt et al., 1981). For a given window $i\in\{1,\dots,k\}$, we use the same notation as in the single-window case for the window and kernel parameters with an added index $i$ denoting the window number: $\delta^{(i)}$, $r^{(i)}$, and $\sigma^{(i)}$. The Gaussian SWR model with multiple windows is given as a direct generalization of (1) by
$y_{t}$ $=\sum_{i=1}^{k}\beta^{(i)}\cdot\left({\bm{x}}_{[t]}\ast\bm{\kappa}^{(i)}\right)+\varepsilon_{t},$ (3)
$={\bm{x}}_{[t]}\ast\left(\sum\limits_{i=1}^{k}\beta^{(i)}\bm{\kappa}^{(i)}\right)+\varepsilon_{t}$ (4)
where the regression parameters $\beta^{(1)},\dots,\beta^{(k)}\geq 0$ act as window weights. In summary, the Gaussian SWR model with $k$ windows contains the following parameters
  • $k$ regression parameters $\bm{\beta}=(\beta^{(1)},\dots,\beta^{(k)}$), $\beta^{(i)}\in\mathbb{R}^{+}$,
  • $k$ lag parameters $\bm{\delta}=(\delta^{(1)},\dots,\delta^{(k)}$), $\delta^{(i)}\in\mathbb{R}^{+}$, and
  • $k$ kernel standard deviation (size) parameters $\bm{\sigma}=(\sigma^{(1)},\dots,\sigma^{(k)}$), $\sigma^{(i)}\in\mathbb{R}^{+}$.
The concept of a Gaussian SWR model with multiple windows is illustrated in Fig. 4: a target value $y_{t}$ is predicted by multiple windows accumulating information in ${\bm{x}}_{[t]}$ over different time intervals. As depicted in Fig. 5, the combined kernel represents a linear combination of the single window kernels $\bm{\kappa}^{(i)}$ for $i=1,\dots,k$. Under the assumption that each window has a distinct centre point $\delta^{(i)}$ (otherwise, identifiability of the parameters might not be given), without loss of generality we can order the location parameters such that $\delta^{(1)}<\delta^{(2)}<\dots<\delta^{(k)}$.
@@
Figure 4: Illustration of the Sliding Windows Regression model with $k=2$ windows, which predicts $y_{70}$ based on $x_{[70]}$. The window location parameters are $\delta^{(1)}=20$ and $\delta^{(2)}=40$. On the time axis, the window $[40,60]$ is centred around $s-\delta^{(1)}=50$ and covers $r^{(1)}=10$ time points on each side of $\delta^{(1)}=50$, while window $[10,50]$ is centred around $s-\delta^{(2)}=30$ and covers $r^{(2)}=20$ time points on each side.
@@ 2.2.1 Parameter interpretation
The number of windows $k$ models the number of underlying, separable flow paths. Since single-window kernels $\bm{\kappa}$ are normalized, regression parameters (window coefficients) $\beta^{(1)},\dots,\beta^{(k)}$ act as weights among the kernels. For practical evaluations, they may be converted into proportions representing the relative impacts of the underlying flow paths.
For a window that is essentially untruncated, the location parameter $\delta^{(i)}$ and the width $\sigma^{(i)}$ of kernel $i$ are loosely related to the mean and the standard deviation, respectively, of the time lag between precipitation and a change in streamflow. In any case $\delta^{(i)}$ is always the window mode and is thus a direct estimate of the typical time delay between rainfall events and the resulting impact on the gauged runoff for a flow path.
@@
Figure 5: Example mixture of window kernels (weights): kernel $\bm{\kappa}^{(1)}$ represents a truncated Gaussian, while kernel $\bm{\kappa}^{(2)}$ has a Gaussian shape. The combined window $\beta^{(1)}\bm{\kappa}^{(1)}+\beta^{(2)}\bm{\kappa}^{(2)}$ is a mixture of Gaussians for the SWR model.
@@ 2.3 Information criteria
$\varepsilon_{t}=y_{t}-{\bm{x}}_{[t]}\ast\left(\sum\limits_{i=1}^{k}\beta^{(i)}\bm{\kappa}^{(i)}\right),$
@@ 2.4 Model training & implementation
    $\bm{D}^{(0)}$ (a vector of starting values for the $\delta^{(i)}$)
    $\bm{S}^{(0)}$ (a vector of starting values for the $\sigma^{(i)}$)
@@
         $\bm{B}^{(1)}\leftarrow(m_{y}/m_{x},\ldots,m_{y}/m_{x})^{\top}$triangleright (for $\beta^{(1)}$)
         $\bm{D}^{(1)}\leftarrow(\bm{D}^{(0)},\bm{D}^{(0)},\ldots)^{\top}$triangleright (for $\delta^{(1)}$)
         $\bm{S}^{(1)}\ \leftarrow(\bm{S}^{(0)},\bm{S}^{(0)},\ldots)^{\top}$triangleright (for $\sigma^{(1)}$)
@@
         Generate an $n\times 1$ vector of starting values for $\beta^{(1)},\ldots,\beta^{(k)}$
@@
                  $\bm{B}^{(j)}\leftarrow(m_{y}/m_{x}/k,\ldots,m_{y}/m_{x}/k)$triangleright (for $\beta^{(j)}$)
@@
         Generate an $n\times 1$ vector of starting values for $\delta^{(k)}$ (note $\bm{D}^{(1)},\ldots,\bm{D}^{(k-1)}$ are unchanged)
@@
         Generate an $n\times 1$ vector of starting values for $\sigma^{(k)}$ (note $\bm{S}^{(1)},\ldots,\bm{S}^{(k-1)}$ are unchanged)
@@
     Generate an $n\times(3k)$ population of starting values with columns for $\beta^{(1)},\ldots,\beta^{(k)},\delta^{(1)},\ldots,\delta^{(k)},\sigma^{(1)},\ldots,\sigma^{(k)}$
@@
We use the GENOUD algorithm (Mebane and Sekhon, 2011) in the R package rgenoud (Mebane et al., 2023) to train a model with a fixed number of windows. GENOUD optimizes a population of candidate solutions over a number of generations via operations analogous to genetic evolution. It also fine tunes the best candidate found at each generation with a quasi-Newton method. Trial and error comparisons with a more standard constrained optimizer (BOBYQA, see Powell, 2009) demonstrated a clear advantage in the optimal log likelihood found, which is important if BIC is to reliably choose the number of windows. As a stopping criterion, the absolute tolerance is set to $10^{-3}$, which is a trivial change in the log likelihood.
@@
In the GENOUD algorithm, we constrain the parameters such that all $\delta^{(i)}>0$ and all $\sigma^{(i)}>1/6$, which allows for truncated as well as non-truncated kernels. In the boundary case of $\delta=0$, a maximum level of truncation is reached for the Gaussian kernel. The constraint $\sigma^{(i)}>1/6$, avoids collapse of the window. Moreover, the minimum window width is hence 1, i.e., ±$\pm 3\times 1/6$, which is commensurate with the limitation of discrete data at unit time steps.
@@
GENOUD requires a population of $n$ starting values for $\beta$, $\delta$, and $\sigma$. Experimental observations showed that this initial population is important. It needs to provide coverage of reasonable values in the parameter space but not be so overly wide as to make the optimization inefficient. Note that the initial ranges can be exceeded later during the GENOUD iterations. The population formed in Alg. 1 has $n$ members in the rows of a matrix ($n=100$ for all results reported) and $3k$ columns for initial values of $\beta^{(1)},\ldots,\beta^{(k)},\delta^{(1)},\ldots,\delta^{(k)},\sigma^{(1)},\ldots,\sigma^{(k)}$, respectively. For $k=1$, $\beta^{(1)}$ is always set to $m_{y}/m_{x}$, where $m_{x}$ and $m_{y}$ are the sample means of $\bm{x}_{T}$ and $\bm{y}_{T}$, respectively, to respect the scaling of the data when a single kernel has total unit weight. The $\delta^{(1)}$ starting values repeat a vector $\bm{D}^{(0)}$; we use $\bm{D}^{(0)}=(1,2,\ldots,25)^{T}$, which will be repeated four times to fill the population of size $n=100$. This range is reasonable for the hydrology application but would be changed in other contexts. Similarly, the $\sigma^{(1)}$ starting values repeat $\bm{S}^{(0)}=(1,5,10,20)^{T}$. When $k>1$ and window $k$ is added, $\beta^{(1)},\ldots,\beta^{(k)}$ are all reduced in magnitude to $m_{y}/m_{x}/k$ to account for $k$ kernels. The starting values for $\delta^{(1)}$ and $\sigma^{(1)}$ are as for the $k=1$ iteration; the $\delta^{(k)}$ and $\sigma^{(k)}$ for the further windows are given by a random permutation of the elements in $\bm{D}^{(0)}$ and $\bm{S}^{(0)}$, respectively. This procedure ensures the starting values for a multi-window iteration have sufficient coverage of the parameter space, while having distinct initial values for each of the $k$ windows.
@@
Note that the suggested model initializations are specific to the hydrological application tackled in this work. A different choice of the hyperparameters may be required if the model was transferred to another domain. The Gaussian SWR model is implemented in R version 4.3.0 (R Core Team, 2022). The implementation is publicly available on GitHub [^1] https://www.github.com/sschrunner/SlidingWindowReg.